OpenStudy (anonymous):

f(x)=x[1 /2 (1−x 2 ) ^-1/ 2 (−2x)]+ sqroot(1−(x^2)) how is it simplified to 1-2x^2=0 ??? please (what are the steps?

OpenStudy (anonymous):

Is the original equation\[f(x) = x\left[\frac{1}{2}(1-x^{2})^{-\frac{1}{2}}(-2x)\right]+ \sqrt{1-x^{2}}\] ? If so, it is written in the Equation input as "f(x) = x\left[\frac{1}{2}(1-x^{2})^{-\frac{1}{2}}(-2x)\right]+ sqrt{1-x^{2}}" without the quotes.